OpenStudy (anonymous):

If x and y are the tens digit and the units digit, respectively, of the product 725,278 x 67,066 what is the value of x+y? a. 12 b. 10 c. 8 d. 6 e. 4

OpenStudy (aaronandyson):

Multiply 725, 278 x 67,066

OpenStudy (aaronandyson):

That is 48641494348 x is tens digit y is units digit Therefore,x = 4 and y = 8,respectively. Now,add x+y Therefore, 4 + 8 Therefore x + y = 4 + 8 = 12

OpenStudy (anonymous):

Hey is there an easier way to do this because I wont be able to use the calculator....